# Particle production and Higgs reheating

### v1

@@ 1 Introcution
The inflationary paradigm [1, 2, 3, 4, 5, 6] provides a compelling explanation for the observed homogeneity, isotropy, and near scale invariance of the cosmic microwave background (CMB) [7, 8, 9]. Yet, the details on how the Universe transitions from the cold, vacuum dominated inflationary state [1, 2, 3, 4, 5, 6] to the hot, radiation dominated era of the Standard Model (SM) [10, 11, 12] remains an open question. This transition, known as reheating [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], governs the initial conditions for all subsequent cosmic evolution and links inflationary dynamics to particle physics.
@@
In the standard picture, reheating proceeds through perturbative decays or parametric resonance of the inflaton into lighter degrees of freedom [16, 17, 18, 19, 20, 21, 22]. However, in many theoretically motivated models; particularly those arising from supergravity [23], string theory [24, 25, 26, 27, 28, 29], or grand unification [30, 31, 32]; the inflaton couples only gravitationally to other fields and, explicit couplings are either suppressed or absent [33, 29, 28, 27, 26, 25, 24]. In such cases, energy transfer to the SM sector must occur through gravitational or geometric effects, a mechanism collectively referred to as gravitational reheating [34, 35]. The study of this process is especially relevant in minimal models of the early Universe, where simplicity and predictive power are preserved by avoiding arbitrary couplings [6, 4, 36, 37, 38].
An intriguing realization of this scenario is provided by Higgs inflation [39, 40, 41, 42, 43], in which the Standard Model Higgs doublet plays the role of the inflaton [39, 36, 44] through a non-minimal coupling to the Ricci scalar. In this work, we revisit reheating such Higgs like inflationary models from a purely geometric standpoint [34, 35, 45, 46, 47, 48]. We consider an inflaton with a general non-minimal coupling to curvature and analyze its gravitational interaction with a set of spectator scalar fields $\chi_{n}$.
In the second part of the paper, we develop the formalism of gravitational particle production [49, 50, 51] using the Bogoliubov transformation of the spectator field modes. By solving the mode equation in a Friedmann Lemaîre Robertson Walker (FLRW) background [52, 53, 54, 55], we determine the occupation numbers, comoving number densities, and the out of equilibrium corrections induced by finite decay rates [35]. This framework enables a quantitative connection between the curvature induced creation of particles and the subsequent thermalization of the universe.
@@ 2 Higgs Reheating
In conventional scenarios, reheating occurs because the inflaton is explicitly coupled to lighter fields through interaction terms such as $g\phi^{2}\chi^{2}$ or Yukawa type couplings [34]. As the inflaton oscillates around the minimum of its potential, these couplings allow perturbative decays or resonant preheating, efficiently producing quanta of other fields [16, 17, 18, 19, 20, 21, 22]. However, such mechanisms depend on model specific choices of coupling constants whose magnitudes are often ad hoc. In many ultraviolet motivated models such as those emerging from supergravity [23], string theory [24, 25, 26, 27, 28, 29], or grand unified constructions [30, 31, 32], the inflaton may interact only gravitationally with the visible sector. In this case, particle creation must proceed through the geometry of spacetime itself. We refer to this minimal, coupling free scenario as geometric reheating [34, 35].
@@
We adopt $\xi_{\nu}$ are equal to $\xi$ (we shall restrict our considerations to positive values of $\xi$ in our numerical evaluations, so as to keep all $\xi_{\nu}$’s negative [^2] In the convention $R=+6(\dot{H}+2H^{2})$, this translates to having $\xi$ negative and having all $\xi_{\nu}$’s positive [^3] See the thesis of Brain R. Greene [56] to understand the phenomenological consequences of assuming $\xi_{\nu}$ to be positive ) as a benchmark case corresponding to symmetric but opposite sign curvature couplings between the inflaton and matter sectors. In more general EFT settings, $\xi_{\nu}$ may differ from $\xi$, a case we leave for future work.
@@ 3 Gravitational Particle Production
When inflation ends, the rapid change in the cosmic expansion rate makes spacetime itself a source of particle creation. Even in the absence of direct couplings between the inflaton and other fields, the time dependent background curvature can excite quantum fluctuations and populate real particles [39]. This process, known as gravitational particle production, is a generic consequence of quantum field theory in curved spacetime [50, 51].
@@
In fig.4(a) we observe that for small $\xi$ ($\xi=0.02$), the curvature oscillations have modest amplitude, leading to particle production localized around the modes that most strongly violate the adiabaticity condition. This produces a characteristic peaked spectrum of $n_{k}$. While fig.4(b) shows that for large $\xi$ ($\xi=1$), curvature oscillations are much stronger, broadening the range of modes affected. The adiabaticity condition is violated for nearly all sub-horizon modes, producing a nearly linear rise of $n_{k}$. Thus, the shape of the spectrum reflects how efficiently curvature oscillations couple to different comoving scales.
@@
In physical terms, gravitational particle production acts as a quantum complement to the classical reheating process discussed earlier. Whereas the geometric reheating described in Sec.2 converts inflaton energy into radiation through macroscopic curvature couplings, the mechanism here originates from microscopic quantum effects associated with a time dependent spacetime metric. Together, the two processes determine how efficiently the Universe can reheat when direct interactions between the inflaton and matter are absent. The combined effect sets the initial radiation temperature and, consequently, influences subsequent phenomena such as baryogenesis, dark matter production, and the generation of relic gravitational waves.
@@ 4 Conclusions
In parallel, we analyzed gravitational particle production [39, 50, 51], a purely quantum process by which fluctuations of spectator fields are excited due to the time dependent background curvature. Using the Bogoliubov formalism, we expressed the number density of produced quanta in terms of the mode functions and demonstrated how deviations from adiabatic evolution translate into particle creation. Furthermore, we encountered the interesting possibility where in the spectator particles can possible be interpreted as LFDM candidates [64, 65], effectively behaving as radiation in the radiation dominated phase but gradually transitioning into behaving as dark matter as the matter-radiation equality approaches. Moreover, extensions to include out of equilibrium effects show that decays or scatterings of these particles can further influence the total reheating efficiency [78, 79].
@@ Acknowledgments
The work of M. K. was performed in Southern Federal University with financial support of grant of Russian Science Foundation № 25-07-IF. We are grateful to Zhong and Ishaan Swamy for their valuable discussions and constant encouragement throughout this work. Finally, we are profoundly grateful to Prof. Pankaj Joshi, Meet Vyas and peers at ICSC for their mentorship, advice, and continuous support.

### v2

@@ 1 Introcution
The inflationary paradigm [52, 65, 3, 66, 72, 81] provides a compelling explanation for the observed homogeneity, isotropy, and near scale invariance of the cosmic microwave background (CMB) [30, 56, 16]. Yet, the details on how the Universe transitions from the cold, vacuum dominated inflationary state [52, 65, 3, 66, 72, 81] to the hot, radiation dominated era of the Standard Model (SM) [22, 46, 41, 48] remains an open question. This transition, known as reheating [62, 28, 36, 60, 61, 51, 50, 4, 70, 12], governs the initial conditions for all subsequent cosmic evolution and links inflationary dynamics to particle physics.
@@
In the standard picture, reheating proceeds through perturbative decays or parametric resonance of the inflaton into lighter degrees of freedom [60, 61, 51, 50, 4, 70, 12, 83, 38]. However, in many theoretically motivated models; particularly those arising from supergravity [27], string theory [74, 88, 32, 19, 31, 67], or grand unification [47, 43, 64]; the inflaton couples only gravitationally to other fields and, explicit couplings are either suppressed or absent [63, 67, 31, 19, 32, 88, 74]. In such cases, energy transfer to the SM sector must occur through gravitational or geometric effects, a mechanism collectively referred to as gravitational reheating [11, 54]. The study of this process is especially relevant in minimal models of the early Universe, where simplicity and predictive power are preserved by avoiding arbitrary couplings [81, 66, 23, 58, 40].
An intriguing realization of this scenario is provided by Higgs inflation [29, 76, 53, 77, 45], in which the Standard Model Higgs doublet plays the role of the inflaton [29, 23, 80] through a non-minimal coupling to the Ricci scalar. In this work, we revisit reheating such Higgs like inflationary models from a purely geometric standpoint [11, 54, 68, 55, 82, 39]. We consider an inflaton with a general non-minimal coupling to curvature and analyze its gravitational interaction with a set of spectator scalar fields $\chi_{n}$.  [^1] See [10] for the impact of scalar fluctuations ($\chi$) non-minimally coupled to gravity, $\xi\chi^{2}R$, as a potential source of secondary gravitational waves (SGWs).
In the second part of the paper, we develop the formalism of gravitational particle production using the Bogoliubov transformation of the spectator field modes [42, 24, 44, 9, 8]. By solving the mode equation in a Friedmann Lemaîre Robertson Walker (FLRW) background [87, 14, 73, 37], we determine the occupation numbers, comoving number densities, and the out of equilibrium corrections induced by finite decay rates [54]. This framework enables a quantitative connection between the curvature induced creation of particles and the subsequent thermalization of the universe.
@@ 2 Higgs Reheating
In conventional scenarios, reheating occurs because the inflaton is explicitly coupled to lighter fields through interaction terms such as $g\phi^{2}\chi^{2}$ or Yukawa type couplings [11]. As the inflaton oscillates around the minimum of its potential, these couplings allow perturbative decays or resonant preheating, efficiently producing quanta of other fields [60, 61, 51, 50, 4, 70, 12, 83, 38]. However, such mechanisms depend on model specific choices of coupling constants whose magnitudes are often ad hoc. In many ultraviolet motivated models such as those emerging from supergravity [27], string theory [74, 88, 32, 19, 31, 67], or grand unified constructions [47, 43, 64], the inflaton may interact only gravitationally with the visible sector. In this case, particle creation must proceed through the geometry of spacetime itself. We refer to this minimal, coupling free scenario as geometric reheating [11, 54].
@@
We adopt $\xi_{\nu}$ are equal to $\xi$ (we shall restrict our considerations to positive values of $\xi$ in our numerical evaluations, so as to keep all $\xi_{\nu}$’s negative [^3] In the convention $R=+6(\dot{H}+2H^{2})$, this translates to having $\xi$ negative and having all $\xi_{\nu}$’s positive. To better understand the phenomenological consequences of assuming $\xi_{\nu}$ to be positive, see the thesis of Brain R. Greene [49].) as a benchmark case corresponding to symmetric but opposite sign curvature couplings between the inflaton and matter sectors. In more general EFT settings, $\xi_{\nu}$ may differ from $\xi$, a case we leave for future work.
@@ 3 Gravitational Particle Production
When inflation ends, the rapid change in the cosmic expansion rate makes spacetime itself a source of particle creation [6]. Even in the absence of direct couplings between the inflaton and other fields, the time dependent background curvature can excite quantum fluctuations and populate real particles [29]. This process, known as gravitational particle production, is a generic consequence of quantum field theory in curved spacetime [24, 44].
@@
In fig.4(a) we observe that for small $\xi$ ($\xi=0.02$), the curvature oscillations have modest amplitude, leading to particle production localized around the modes that most strongly violate the adiabaticity condition. This produces a characteristic peaked spectrum of $n_{k}$. While fig.4(b) shows that for large $\xi$ ($\xi=1$), curvature oscillations are much stronger, broadening the range of modes affected. The adiabaticity condition is violated for nearly all sub-horizon modes, producing a nearly linear rise of $n_{k}$. Thus, the shape of the spectrum reflects how efficiently curvature oscillations couple to different co-moving scales.
@@
In physical terms, gravitational particle production acts as a quantum complement to the classical reheating process discussed earlier. Whereas the geometric reheating described in Sec.2 converts inflaton energy into radiation through macroscopic curvature couplings, the mechanism here originates from microscopic quantum effects associated with a time dependent spacetime metric. Together, the two processes determine how efficiently the Universe can reheat when direct interactions between the inflaton and matter are absent. The combined effect sets the initial radiation temperature and, consequently, influences subsequent phenomena such as baryogenesis, dark matter production [^6] Interested reader can see [5] for the case that if cosmic inflation was driven by an electrically neutral scalar field stable on cosmological time scales then the field necessarily constitutes all or part of dark matter (DM) which is also an important line of investigation to take into account the non-perturbative nature of particle production during reheating., and the generation of relic gravitational waves.
@@ 4 Conclusions
In parallel, we analyzed gravitational particle production [29, 24, 44], a purely quantum process by which fluctuations of spectator fields are excited due to the time dependent background curvature. Using the Bogoliubov formalism, we expressed the number density of produced quanta in terms of the mode functions and demonstrated how deviations from adiabatic evolution translate into particle creation. Furthermore, we encountered the interesting possibility where in the spectator particles can possible be interpreted as LFDM candidates [78, 85], effectively behaving as radiation in the radiation dominated phase but gradually transitioning into behaving as dark matter as the matter-radiation equality approaches. Moreover, extensions to include out of equilibrium effects show that decays or scatterings of these particles can further influence the total reheating efficiency [18, 25]. Also an interesting case arises when we use classical lattice simulations in 3+1 dimensions to study the interplay between the resonant production of particles during preheating and the subsequent decay of these into a set of secondary species [75]. This lattice-based extension to the non-linear regime when applied to this case can reveal some other important aspects which are worthwhile exploring in the future work.
@@ Acknowledgments
The work of M. K. was performed in Southern Federal University with financial support of grant of Russian Science Foundation № 25-07-IF. The work of O.T. was supported in part by the Vanderbilt Discovery Doctoral Fellowship. We are grateful to Robert Brandenberger, Javier Rubio, Ayan Chakraborty, Jiahang Zhong and Ishan Swamy for their valuable discussions and comments. Finally, we are also grateful to Prof. Pankaj Joshi, Meet Vyas, Aum Trivedi and peers at ICSC for their mentorship, advice, and continuous support.